OpenStudy (timaashorty):

Anybody Can help with this? I just started learning about quadratic functions. Please also give details, I'm not just looking for the answer, but also I want to understand how. •Write a real-world problem that can be modeled using a quadratic equation.

OpenStudy (anonymous):

for this all u need to know is the quadratic formula basically

OpenStudy (anonymous):

ur in alegbra i take it?

OpenStudy (timaashorty):

f(x)=a^2+b+c ?

OpenStudy (anonymous):

listen timaasshorty

OpenStudy (anonymous):

go to mathway.com, with that website u can like get an A in alegbra

OpenStudy (anonymous):

it can explain how to solve problems too, if u pay, or u can create facebook accounts over and over and get 1 day free trial

OpenStudy (timaashorty):

thanks for letting me know, I use the website, too. But I like this one also. Was I right about the formula?

OpenStudy (anonymous):

mathway.com is like a calculator

OpenStudy (anonymous):

smart math computer something like that and as for the formula

OpenStudy (anonymous):

google the formula lol

OpenStudy (timaashorty):

I know but I'm not looking for a calculator.

OpenStudy (anonymous):

ax2+bx+c=0

OpenStudy (anonymous):

it can explain how to solve problems step by step too

OpenStudy (timaashorty):

I'm looking for help with this: Write a real-world problem that can be modeled using a quadratic equation

OpenStudy (anonymous):

ax^2+bx+c=0

OpenStudy (anonymous):

real world problem? its probably a word problem it wants u to make

jimthompson5910 (jim_thompson5910):

quadratic functions model parabolas something like an arch is a parabola you could also model the time it takes an object to hit the ground if you threw it up in the air using a quadratic function

OpenStudy (anonymous):

real word quadratic problem: http://www.mathsisfun.com/algebra/quadratic-equation-real-world.html

OpenStudy (anonymous):

When an object is thrown or projected, it follows a curved path called a trajectory. If we neglect air resistance, and assume that the object is fairly dense, and that the object is close to the earth's surface, the path could be described by: H = vt + h -.5at^2. H = the height of the object above its initial position (usually taken as the ground) after some elapsed time v = the initial velocity of the object t = the elapsed time h = the initial height of the object above ground a = the acceleration due to gravity: approximately 32 feet per second squared or 9.8 meters per second squared.

OpenStudy (timaashorty):

Lol I know that sam, but I'm confused whether there are specific numbers I need to apply to a qaudatric formula. or I could insert any numbers ?

OpenStudy (anonymous):

quadratic formula: y = ax^2+bx+c.

OpenStudy (anonymous):

Very simple is the surface of the room. What is the surface of square shaped room of height 3meters whit floor, ceiling and walls walls are 3*m*(4*x) floor and ceiling is 2*x^2 together is 2*x^2+12[m]*x and is square function

OpenStudy (anonymous):

most common types as the other guy stated: One of the most common real life quadratic functions deals with projectile motion - that formula is h(t)=-16t^2+vt+c, where v represents the initial velocity and c represents the initial height.... so think of throwing a ball up in the air if it is thrown from a distance of 3 m above the ground with a velocity of say 10 the equation that would represent the height of the ball at any given time after release would be h(t)=-16t^2+10t+3....from there any t (representing the number of seconds since release of ball) it will give the height of the ball...you could also find the maximum height of the ball and how long it takes to get there (vertex) and when the ball will hit the ground (use quadratic formula and solve when h(t) is 0).... that can get you started, maybe?

OpenStudy (anonymous):

ur welcome

OpenStudy (timaashorty):

thanks for all that information, I needed it, y=a(x+2)^2+3 That would be a quadratic equation?

OpenStudy (anonymous):

yes it would i believe

OpenStudy (timaashorty):

I just noticed you just gavee an example . Thanks so much (: So that's all I needed!